40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Fri Apr 30 12:07:25 2021, aaron, DailyProgress, Laser, marconi locking 252D9563-0122-418B-9B97-29D44FEF83E9.jpeg
    Reply  Fri Apr 30 15:15:33 2021, rana, DailyProgress, Laser, marconi locking 
       Reply  Mon May 3 23:28:56 2021, rana, DailyProgress, Laser, delay line frequency discriminator DFD.jpg
          Reply  Wed May 5 11:54:19 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 9x
             Reply  Wed May 5 22:21:17 2021, rana, DailyProgress, General, delay line frequency discriminator 
             Reply  Thu May 6 10:48:40 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 3EFFFA27-AAF6-434F-A1AB-FE7C657A464C.jpegDL_3CH.pdf
                Reply  Tue May 11 14:37:48 2021, aaron, DailyProgress, Laser, delay line frequency discriminator, current noise measurement 
                   Reply  Wed May 12 15:35:09 2021, aaron, DailyProgress, Laser, delay line in a box 69A8E6A2-7F85-464E-BDAA-8C6BBD2910B5.jpeg1D05A31E-7DBC-448B-A2F4-A8805EFAC393.jpeg25C2DD1E-7776-491E-BBCE-5F2D1EA935E3.jpeg
                   Reply  Wed May 12 16:52:08 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 3878D102-C5BB-46AE-B90B-540CFDE3100B.jpegSPSR785_NX_12-05-2021_173959.pdfSPSR785_SX_12-05-2021_183434.pdfSPSR785_NS_12-05-2021_191513.pdf
                      Reply  Thu May 13 16:31:40 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 6x
                         Reply  Thu May 13 19:09:16 2021, rana, DailyProgress, Laser, delay line frequency discriminator 
                         Reply  Fri May 14 18:43:25 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 994ECC7F-C5FE-4E7B-8F19-B1C118DCE6CE.png
                            Reply  Thu May 20 12:50:56 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 
                               Reply  Fri May 21 14:15:21 2021, aaron, DailyProgress, Laser, current noise of custom current drivers 5012D147-8B13-478C-BB83-57A71A2B304A.jpeg467DEC59-01B1-40CD-9F4E-4D9D87183A78.jpegABC2A51B-6726-4F9C-AD88-551ED7C90EEF.jpeg
                               Reply  Tue Aug 17 14:04:27 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 3A1E242C-F4F6-4A87-B2E8-E74839BE542C.jpeg
                                  Reply  Wed Aug 18 09:55:13 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 0A96102F-5465-456B-93BF-E2B9D4A7E96C.jpegtrim.2553D656-98A0-442D-9185-4C5252B2B727.MOVtrim.787F04FC-8DED-4501-B67D-FF98EC5D5D56.MOV
                                     Reply  Thu Aug 19 14:21:35 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 007FC2D7-F29B-4636-9CE7-04CBC3B789F7.jpeg1F1E68F0-199F-4E1C-B909-883A521E06CF.jpegScreenshot_from_2021-08-19_17-53-58.png
                                        Reply  Fri Aug 20 11:03:44 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 30A43669-5DFF-400A-BEB4-7EF71BBAE1A6.jpegA3BF5DA5-32E5-42A5-A1D8-9270C3D88518.pngScreenshot_from_2021-08-20_16-11-26.pngScreenshot_from_2021-08-20_16-20-48.png
                                           Reply  Mon Aug 23 13:38:57 2021, aaron, DailyProgress, Laser, delay line frequency discriminator Screenshot_from_2021-08-23_16-42-48.png
                                              Reply  Tue Aug 24 16:44:39 2021, aaron, DailyProgress, Laser, delay line frequency discriminator beats_regions.pdfTeraxion_noise.pdfEW_timeseries.pdf
                                                 Reply  Thu Aug 26 11:58:50 2021, aaron, DailyProgress, Laser, delay line frequency discriminator 9x
                                                    Reply  Fri Aug 27 15:27:38 2021, rana, DailyProgress, Laser, delay line frequency discriminator 
                                                    Reply  Tue Aug 31 15:14:28 2021, aaron, DailyProgress, Laser, quieting the Rio W and Teraxion beat notes, but not much help beats_regions_base2.pdfbeats_regions_base2.pdfTeraxion_noise_base2.jpgDLFD_foam.pdf
                                                       Reply  Wed Sep 1 12:20:18 2021, rana, DailyProgress, Laser, quieting the Rio W and Teraxion beat notes, but not much help 
                                     Reply  Thu Aug 19 16:32:47 2021, rana, DailyProgress, Laser, delay line frequency discriminator 
       Reply  Tue May 4 10:50:22 2021, aaron, DailyProgress, Laser, marconi locking 6DC5CB12-BCDF-4550-AFB8-42D3DC354AED.jpegSPSR785_04-05-2021_152326.pdf
          Reply  Wed May 5 00:27:10 2021, rana, DailyProgress, Laser, marconi locking 
Message ID: 2737     Entry time: Tue May 11 14:37:48 2021     In reply to: 2730     Reply to this: 2738   2739
Author: aaron 
Type: DailyProgress 
Category: Laser 
Subject: delay line frequency discriminator, current noise measurement 

I'm repeating this measurement, this time recording the current noise of the commercial driver before each measurement for reference. Will update tomorrow with additional measurements, plots.

Delay line frequency discriminator procedure

  1. Set up the delay line frequency discriminator as before (see attachment 1)
  2. Find a low noise operating point (see below) for the Rio laser(s)
  3. Calibrate the delay line by tuning to the dark and bright fringes around 80 MHz and recording the DC voltage of the LF (after whatever amplification is applied)
  4. Measure the current noise of the commercial driver as below [could also be done completely independent of these steps]
  5. Tune the beat note frequency to the half fringe, and record the beat note spectrum.
  6. Repeat for each pair of beat notes.
  7. When done, turn off RF amplifier (first disconnect input, then turn off DC power) and SR560, and remember to plug in the SR560 power line.

The measurement parameters for delay line frequency discriminator are:

  • Pickoff RF from 1611 to Moku (50 Ohm AC coupled spectrum analyzer) using
  • RF amplifier ZHL-1A (powered by +24 V from GPS3030D) to splitter
  • Level 7 mixer ZFM-3-S+ (0.04-400 MHz)
  • LF from mixer through SMA-T with one end terminated at 50 Ohms, the other to SLP-1.9+ then SR560 channel A
  • SR560 settings: battery powered, DC coupled, 30 kHz LP (6 dB/oct), low noise mode, G=10
laser pair Beat note power Rio N laser TEC setpoint lower null frequency lower null DC voltage lower null LD current upper null frequency upper null DC voltage Upper null LD current voltage at half fringe frequency at half fringe RMS voltage at half fringe Hz/V calibration start time of measurement drift during measurement notes
Rio N x Teraxion 440 uV 9.548 kOhm 63.5 MHz .18 V 95.5 mA 101.1 MHz -3.28 V 96.2 mA -.95 V 82 MHz 35 mV   Tue May 11 17:26:42 2021   1dB attenuation on LO

 

Finding a 'low noise' operating point

According to the datasheet (PSOMA wiki 'documents' page), the lasers come with a recommended temperature setpoint (T_set on the datasheet). This setpoint may either lie on the upper hysteresis branch, or lie outside of the hysteresis region... but according to our datasheet, our lasers' T_set lie in the upper hysteresis branch. If we observe spectral distortions, this indicates the temperature is outside the recommended range or operating on the lower arm of the hysteresis. To achieve a low noise operating point

  1. Ensure the laser is properly mounted on the PCB with drivers and control loops connected (but off)
  2. activate the thermal control loop, and let the internal temperature stabilize
  3. Apply the laser bias current, ramping the current at 10 mA/sec up to I_bias
  4. For the delay line measurement, scan the laser current until there is a beat note near 80 MHz
    • The 1611 may be saturating. If harmonics of the main beat note are visible,
      • iteratively reduce the laser current and adjust the TEC setpoint while keeping the beat note near 80 MHz
      • Since the laser power depends more strongly on current, this will reduce the beat note power. Do so until no harmonics are visible on the spectrum analyzer.
      • Empirically, I found this to be true when the total power on the photodiode was under 500 uW

Notes about the lasers...

  • If it is required to set the bias current and TEC current simultaneously, first tune the laser to several degrees below the set point in the hysteresis free region, then slowly increase the TEC temperature to the setpoint
  • Ensure that the TEC control loop is restricted from entering the mode hop region, since this will put us on the lower hysteresis curve
  • For the North laser (SN: 104978), T_set is 25 C, corresponding to about 9.9-10.1 kOhm. Operated at the recommended I_bias of 150 mA, I observed the next mode hop at 8.971 kOhm, so the TEC control should be kept above that level (operated at lower bias currents, the hop has been observed at lower values of TEC resistance).
  • For the South laser (SN: 104987), T_set is 23 C, corresponding to about 10.8-11.04 kOhm.

 Measuring current noise

  1. Turn off the laser by
    1. Ramp down then turn off LD current
    2. Turn off TEC
    3. Turn off driver box
  2. Remove DB9 at the driver box, and add a DB9 breakout between the driver and the cable going to the laser.
    • The voltage across the diode is between pins 3 and 7
  3. Send the voltage across pins 3 and 7 to a floating-input DC coupled SR560, and the output of the SR560 to an oscilloscope
  4. Turn on the laser by
    1. Turn on driver box
    2. Turn on TEC
    3. Turn on LD current by
      1. With the current off, turn the current setpoint to 0
      2. Turn on the LD current
      3. Ramp the LD current by 10 mA/s up to the desired setpoint
  5. Note the voltage across the diode at 2-3 different values of bias current near the operating point. This is the voltage to current calibration.
  6. Turn off the laser as above
  7. Switch the SR560 to AC coupled mode by
    1. Disconnect the BNC from the SR560 input
    2. switch the preamplifier to AC coupled mode
    3. Change the gain to G=100
    4. reconnect the BNC to SR560 input.
  8. Turn on the laser as above
  9. Record the output of the SR560 with a spectrum analyzer. Note the settings of the SR560, and use that and the calibration above to convert the spectrum from V/rtHz to A/rtHz of the current driver.
    • The delay line frequency discriminator calibration can then be used to convert from A/rtHz to Hz/rtHz
  10. Return the system to its original state by
    1. Turn off laser as above
    2. Remove DB9 breakout and reattach DB9 cable directly to LD driver
Laser TEC setpoint diode setpoint (I_0) Voltage at I_0 diode current (I_1) Voltage at I_1 diode current (I_2) Voltage at I_2 R_sense

Rio N

(ITC510)

9.548
 kOhm
 95.9
 mA
-1.84 V 100.4 mA -1.89 V 92.5
 mA
-1.82 V 8.98 Ohm
Rio S (ITC502) 9.515 kOhm 122.92 mA -2.15 V 119.05 mA -2.11 V 126.53 mA -2.18 V 9.37 Ohm

Notes on current noise measurements:

  • SR560 settings: G=100, AC coupled, 30 kHz LP with 6 db/oct, battery powered
  • Updated on May 14 with S laser drive (ITC502) current noise. Measured at 123.43 mA.
ELOG V3.1.3-